A glass flask whose volume is 1000 cm3 at a temperature of 1.00 ∘C is completely filled with mercury at the same temperature. When the flask and mercury are warmed together to a temperature of 52.0 ∘C , a volume of 8.05 cm3 of mercury overflows the flask

If the coefficient of volume expansion of mercury is βHg = 1.80×10−4 /K , compute βglass, the coefficient of volume expansion of the glass.
Express your answer in inverse kelvins.
Answer #1

Here ,

for the coefficient of volume expansion of the glass is β

volume flowed = V * delta T * (βHg - β)

8.05 * 10^-3 = 1000 * 10^-3 * (52 - 1) * (1.80 *10^-4 - β)

solving for β

β = 2.215 *10^-5 1/K

the coefficient of volume expansion of the glass is 2.215 *10^-5 1/K